# Function-Correcting Codes for Insertion-Deletion Channels

### v1

@@ Abstract
In coding theory, handling errors that occur when symbols are inserted or deleted from a transmitted message is a long-standing challenge. Optimising redundancy for insertion and deletion channels remains a key open problem, with significant importance for applications in DNA data storage and document exchange. Recently, a new coding framework known as function-correcting codes has been proposed to address the challenge of optimising redundancy while preserving particular functions of the message. This framework has gained attention due to its potential applications in machine learning systems and long-term archival data storage. To address the problem of redundancy optimisation in insertion and deletion channels, we propose a new coding framework called function-correcting codes for insertion and deletion channels.
In this paper, we introduce the notions of function-correcting insertion codes, function-correcting deletion codes, and function-correcting insdel (insertion–deletion) codes, and we demonstrate that these three formulations are equivalent. We then introduce insdel distance matrices and irregular insdel-distance codes, and further derive lower and upper bounds on the optimal redundancy achievable by function-correcting codes for insdel channels. Furthermore, we establish Gilbert-Varshamov and Plotkin-like bounds on the length of irregular insdel-distance codes. By utilising the relation between optimal redundancy and the length of irregular insdel distance codes, we provide another simplified lower bound on optimal redundancy. We subsequently find bounds on optimal redundancy of FCIDCs for various classes of functions, including locally bounded functions, VT syndrome functions, the number-of-runs function, and the maximum-run-length function.
Index Terms:
Function-correcting codes, error-correcting codes, insertion-deletion channels, optimal redundancy.
I Introduction
@@
FCCs have been commonly studied under systematic encoding, which is important in applications such as distributed computing and long-term archival storage, where preserving the original data is essential. Since redundancy is the primary and most crucial factor for FCCs, it becomes crucial to precisely determine optimal redundancy—that is, the minimum number of additional bits required to guarantee correct computation of the intended function. To determine optimal redundancy for both generic and specific functions, the authors of [10132545] introduced the concept of irregular-distance codes, which are closely connected to the redundancy patterns of FCCs. They established equivalence between optimal redundancy of FCCs and shortest achievable length of irregular-distance codes. Using this equivalence, the authors derived redundancy bounds for arbitrary functions and subsequently used these results to obtain tight or near-tight bounds for several important function classes, including locally binary bounded functions, Hamming weight functions, Hamming weight distribution functions, and real-valued functions.
I-A Motivation
@@
The nature of the underlying channel noise fundamentally determines how function values are distorted during transmission. By designing FCCs that are tailored to specific channel models, one can achieve reliable function computation in scenarios where recovering the entire message is either unnecessary or prohibitively costly in redundancy. Moreover, such channel-specific constructions provide deeper insight into the way different error processes affect function evaluation. Motivated by this observation, we focus on selecting suitable channels and associated functions to construct function-correcting codes with improved redundancy. In particular, we consider insertion–deletion(insdel) channels, where errors occur when symbols are inserted into or deleted from a transmitted message. Such errors disrupt the alignment between transmitted and received sequences, creating a form of noise that traditional error-correcting codes designed primarily for handling substitution errors are not well-suited to handle. As these types of errors frequently arise in biological systems, modern storage devices, and asynchronous communication settings, the study of insdel channels has become increasingly important in developing robust and efficient coding techniques. The study of insertion–deletion errors began in the 1960s [varshamov1965codes, levenshtein1965двоичные, levenshtein1967asymptotically] and [tenengolts1984nonbinary], when V. Levenshtein introduced the edit distance, defining the minimum number of insertions, deletions, or substitutions required to transform one string into another. This concept established the mathematical basis for analyzing synchronization errors, although early work focused mainly on error detection and correction rather than determining full channel capacities. Later, practical coding schemes for handling such errors began to emerge, most notably the Varshamov–Tenengolts (VT) codes, which were originally introduced in [varshamov1965codes] and expanded over subsequent decades, enabling reliable single-deletion correction. During this period, researchers also explored the use of synchronizing sequences to help maintain alignment between transmitted and received data, marking a significant step toward systematically managing insertion-deletion errors. Consequently, the development of function-correcting codes under the insdel metric represents both a substantial theoretical milestone and a vital practical capability for next-generation communication and computation systems. By offering resilience against synchronization errors—most notably insertions and deletions—these codes broaden the scope of traditional error-control methods beyond the limitations of the Hamming metric, enabling more reliable transmission, storage, and processing of information in emerging high-density, high-throughput environments. The nature of the underlying noise fundamentally influences how function values are distorted during transmission. A function that is robust to substitution errors may behave unpredictably under insertions, deletions, erasures, or symbol-dependent noise, and the structural interactions between the function and the channel can vary dramatically. Developing FCCs tailored to specific channels enables reliable computation in settings where full message recovery is either unnecessary or prohibitively expensive, while also revealing deeper insights into how functions transform under diverse error processes. Such channel-aware FCC frameworks are essential for modern applications ranging from asynchronous communication and DNA storage to distributed sensing and edge computing, where computation must remain accurate despite highly nonuniform and context-dependent noise.
I-B Related Work
@@
Function-correcting codes (FCCs) were introduced by Lenz et al. in [10132545] as a coding framework designed to recover the value of a target function of the data rather than the data itself, resulting in substantially lower redundancy compared to traditional error-correcting codes. Their work formalized FCCs, established an equivalence with irregular-distance codes, and provided general upper and lower bounds demonstrating how the structure of the protected function influences the required redundancy. Subsequent developments, such as [Xia2023], revisited the model and extended the framework from the Hamming metric to the symbol-pair metric, deriving several bounds and structural insights. In [Premlal2024], the authors established a lower bound on the redundancy of FCCs for linear functions. Focusing on the Hamming weight and Hamming weight distribution, Ge et al. [Ge2025] produced stronger lower bounds on redundancy and introduced code designs that exactly meet those bounds.
@@
A growing line of research has further generalized FCCs to nonstandard error models. In particular, [Singh2025] extended function-correcting codes to the $b$-symbol read channel, addressing clustered read errors common in modern storage systems. That work introduced irregular $b$-symbol distance codes and derived bounds characterizing the redundancy needed to recover a function value under this metric. The results, supported by a graph-theoretic formulation and illustrative examples, show that FCCs can achieve substantially lower redundancy than traditional $b$-symbol error-correcting codes when only the function output must be protected.
@@
Further extensions were provided in [Verma2025FCCL], which studied function-correcting codes for locally bounded $b$-symbol functions over $b$-symbol read channels. Recently, function correction has also been explored in the Lee-metric setting [rajan2025plotkin, Verma2025FCCL]. More recent work, such as [rajput2025function], investigates hybrid protection models in which both the function value and selected parts of the data must remain resilient to errors, highlighting the versatility of FCCs as a targeted and efficient alternative to full-message protection.
I-C Contributions
@@
I-D Organization
The rest of the paper is organized as follows. Section II provides a comprehensive review of the foundational concepts used throughout the paper. We begin by recalling the basic definitions and properties of insertion–deletion (insdel) codes, including the insdel metric and several equivalent formulations that are relevant to our analysis. In Section III, we introduce the framework of function-correcting deletion codes, function-correcting insertion codes, and function-correcting insertion–deletion (insdel) codes, which are designed to recover the value of a target function in the presence of deletion, insertion, or combined insertion–deletion errors. We then present and discuss irregular insdel-distance codes, highlighting their significance in characterizing optimal redundancy and in deriving key performance limits. Building on this framework, we derive lower and upper bounds on the optimal redundancy achievable by function-correcting codes for insdel channels. In addition, we establish Gilbert–Varshamov-type and Plotkin-like bounds on the lengths of irregular insdel-distance codes, which play a central role in analysing and understanding the redundancy of such codes. In Section IV, we apply the general theoretical results to a specific class of functions, the VT syndrome function, which plays a central role in the insertion–deletion coding. We analyze how the structure of the VT syndrome function interacts with insdel errors and derive both upper and lower bounds on the optimal redundancy required for function correction in this setting. We illustrate the general framework using three concrete classes of functions: the number-of-runs function (Section V), the maximum-run-length function (Section VI), and locally bounded functions (Section VII). For each class, we derive the corresponding redundancy values as well as the associated upper and lower bounds. Section VIII concludes the paper.
I-D1 Notations
@@
II Preliminaries
In this section, we review the basic concepts and definitions that will be used throughout the paper. We begin with a brief discussion of insertion, deletion, and insdel codes, which are designed to enable reliable communication over channels that produce synchronization errors. We then introduce the fundamental notions associated with irregular-distance codes, a class of codes characterized by non-uniform or structure-dependent distance measures. Throughout the paper, we focus on binary insdel codes, where the alphabet is $\mathbb{F}_{2}=\{0,1\}$.
II-A Insdel code
In this paper, we deal with a channel model in which errors occur in the form of deletions, insertions, or a combination of both. Codes designed to handle such types of errors were first introduced by Levenshtein, Varshamov, and Tenengolts in the 1960s [varshamov1965codes, levenshtein1965двоичные, levenshtein1967asymptotically], and [tenengolts1984nonbinary]. In this section, we provide a brief overview of these classical codes and summarize the main results associated with them.
Let $x$ be a binary vector of length $k$ that, when transmitted through a deletion channel, may lose some of its bits. If the channel deletes $t$ bits, the resulting binary vector $y$ has length ktk-t. The vector $y$ is referred to as a subsequence of $x$. A subsequence of a vector is obtained by selecting a subset of its symbols and aligning them in their original order, without any rearrangement. Formally, the subsequence relation is defined as follows:
Definition 1 (Subsequence).
@@ II-A Insdel code
Similarly, when $x$ is transmitted through an insertion channel, additional bits may be inserted at arbitrary positions, resulting in a supersequence of $x$. In the deletion case, the original vector $x$ is referred to as a supersequence of the received vector $y$. In the deletion case, the original vector $x$ is referred to as a supersequence of the received vector $y$. The set of all subsequence of $x$ obtained by deletion of $t$ bits is denoted by $D_{t}(x)$ and set of supersequences obtained by insertion of $t$ bits is denoted by $I_{t}(x)$.
@@
Definition 2 (Deletion Correcting Codes).
[levenshtein1965двоичные] A $t$-deletion-correcting code $\mathcal{C}$ of length $n$ is a subset of binary vector space $\mathbb{F}_{2}^{n}$ that holds the following property for all vectors $x,y\in\mathcal{C}$.
$D_{t}(x)\cap D_{t}(y)=\phi.$
Definition 3 (Insertion Correcting Codes).
[levenshtein1965двоичные] A $t$-insertion correcting code $\mathcal{C}$ of length $n$ is a subset of binary vector space $\mathbb{F}_{2}^{n}$ that holds the following property for all vectors $x,y\in\mathcal{C}$.
$I_{t}(x)\cap I_{t}(y)=\phi.$
@@
Definition 4 (Insdel Metric).
[levenshtein1965двоичные] Let $x$ and $y$ be binary sequences of length $k$. The insdel metric $d_{ID}(x,y)$ between $x$ and $y$ is defined as the minimum number of insertions and deletions required to transform $x$ into $y$.
@@
Definition 5.
(Insdel Code)[do2021explicit] A subset of $\mathbb{F}^{k}_{2}$ is defined to be an insdel code $\mathcal{C}$ whose minimum distance is given by:
@@
Remark 1.
[levenshtein1965двоичные] Any code that can correct $s$ deletions (or equivalently, any code that can correct $s$ insertions) can also correct $s$ combined deletions and insertions.
@@
Lemma 1.
[liu2022bounds] Let $x,y$ be two binary words of length $k$. If $d_{H}(x,y)$ denotes the hamming distance between $x$ and $y$ then we have:
@@
Lemma 2.
@@
Without loss of generality, assume that $r(x)>r(y)$ and set $\Delta=r(x)-r(y)$. A single insertion into a binary string can change the number of runs by at most $2$. Let us assume $d$ insertion operations are required to increment the number of runs of $y$ by $\Delta$. Because each insertion operation changes the run-count by at most 2, we have
$\Delta\leq 2d\implies d\geq\Big\lceil\frac{\Delta}{2}\Big\rceil$
Now, since the length of $x$ and $y$ is equal, any insertion performed to increase the run-count must be compensated by a deletion in order to maintain length $n$. So the total number of operations required is twice the number of insertions required. Combining with the previous inequality, we get,
$d_{ID}\geq 2\Big\lceil\frac{\Delta}{2}\Big\rceil\geq\Delta.$
If $\Delta$ is odd $2\Big\lceil\frac{\Delta}{2}\Big\rceil=\Delta+1.$
For positive integers $k\geq t\geq 1$ and $x\in\mathbb{F}_{2}^{k}$, denote
$\scriptstyle S_{D}(x,t)\triangleq\{v\in\mathbb{F}_{2}^{k-t}:v\ \text{is obtained from $\scriptstyle x$ by deleting $\scriptstyle t$ symbols}\},$
$\scriptstyle S_{I}(x,t)\triangleq\{v\in\mathbb{F}_{2}^{k+t}:v\ \text{is obtained from $\scriptstyle x$ by inserting $\scriptstyle t$ symbols}\},$
as the deletion sphere and insertion sphere centered at $x$ of radius $t$, respectively. For even number $2\leq d\leq 2k$, denote
@@
as the insdel ball centered at $x$ of radius $t$. Then, we have the following bounds on $|S_{D}(x,t)|$, $|S_{I}(x,t)|$ and $|B_{E}(x,d)|$.
Lemma 3.
@@
  1. 1.
    $\binom{\rho(x)+t-1}{t}\leq|S_{D}(x,t)|\leq\binom{k+t}{t}$, where $\rho(x)$ is the number of runs( A run is a maximal sequence of identical symbols.) in $x$.
  2. 2.
    $|S_{I}(x,t)|\leq\sum_{i=0}^{t}\binom{k+t}{i}$.
  3. 3.
    $|B_{ID}(x,2t)|\leq\sum_{i=0}^{t}|S_{D}(x,i)|\cdot|S_{I}(x,i)|\leq k^{2t}2^{t}\left(1+o_{k}(1)\right)\quad\text{as }k\to\infty.$
Lemma 4.
@@
II-B Function-Correcting Codes(FCC)
Definition 6 ([10132545]).

### v2

@@ Abstract
In coding theory, handling errors that occur when symbols are inserted or deleted from a transmitted message is a long-standing challenge. Optimising redundancy for insertion and deletion channels remains a key open problem, with significant importance for applications in DNA data storage and document exchange. Recently, a new coding framework known as function-correcting codes has been proposed to address the challenge of optimising redundancy while preserving particular functions of the message. This framework has gained attention due to its potential applications in machine learning systems and long-term archival data storage. To address the problem of redundancy optimisation in insertion and deletion channels, we propose a new coding framework called function-correcting codes for insertion-deletion channels.
In this paper, we introduce the notions of function-correcting insertion codes, function-correcting deletion codes, and function-correcting insdel (insertion–deletion) codes(FCIDCs), and we demonstrate that these three formulations are equivalent. We then introduce insdel distance matrices and irregular insdel-distance codes, and further derive lower and upper bounds on the optimal redundancy achievable by function-correcting codes for insdel channels. Furthermore, we establish Gilbert-Varshamov and Plotkin-like bounds on the length of irregular insdel-distance codes. By utilising the relation between optimal redundancy and the length of irregular insdel-distance codes, we provide another simplified lower bound on optimal redundancy. We subsequently find bounds on optimal redundancy of FCIDCs for various classes of functions, including locally bounded functions, VT syndrome functions, the number-of-runs function, and the maximum-run-length function.
Keywords: Function-correcting codes, error-correcting codes, insertion-deletion channels, optimal redundancy.
1 Introduction
@@
FCCs have been commonly studied under systematic encoding, which is important in applications such as distributed computing and long-term archival storage, where preserving the original data is essential. Since redundancy is the primary and most crucial factor for FCCs, it becomes crucial to precisely determine optimal redundancy—that is, the minimum number of additional bits required to guarantee correct computation of the intended function. To determine optimal redundancy for both generic and specific functions, the authors of [1] introduced the concept of irregular-distance codes, which are closely connected to the redundancy patterns of FCCs. They established equivalence between the optimal redundancy of FCCs and the shortest achievable length of irregular-distance codes. Using this equivalence, the authors derived redundancy bounds for arbitrary functions and subsequently used these results to obtain tight or near-tight bounds for several important function classes, including locally binary bounded functions, Hamming weight functions, Hamming weight distribution functions, and real-valued functions.
1.1 Motivation
@@
The nature of the underlying channel noise fundamentally determines how function values are distorted during transmission. By designing FCCs that are tailored to specific channel models, one can achieve reliable function computation in scenarios where recovering the entire message is either unnecessary or prohibitively costly in redundancy. Moreover, such channel-specific constructions provide deeper insight into the way different error processes affect function evaluation. Motivated by this observation, we focus on selecting suitable channels and associated functions to construct function-correcting codes with improved redundancy. In particular, we consider insertion–deletion(insdel) channels, where errors occur when symbols are inserted into or deleted from a transmitted message. Such errors disrupt the alignment between transmitted and received sequences, creating a form of noise that traditional error-correcting codes, designed primarily for handling substitution errors, are not well-suited to handle. As these types of errors frequently arise in biological systems, modern storage devices, and asynchronous communication settings, the study of insdel channels has become increasingly important in developing robust and efficient coding techniques. The study of insertion–deletion errors began in the 1960s [2, 3, 4], and [5], when V. Levenshtein introduced the edit distance, defining the minimum number of insertions, deletions, or substitutions required to transform one string into another. This concept established the mathematical basis for analyzing synchronization errors, although early work focused mainly on error detection and correction rather than determining full channel capacities. Later, practical coding schemes for handling such errors began to emerge, most notably the Varshamov–Tenengolts (VT) codes, originally introduced in [2] and subsequently expanded, enabling reliable single-deletion correction. During this period, researchers also explored the use of synchronizing sequences to help maintain alignment between transmitted and received data, marking a significant step toward systematically managing insertion-deletion errors. Consequently, the development of function-correcting codes under the insdel metric represents both a substantial theoretical milestone and a vital practical capability for next-generation communication and computation systems. These codes extend the reach of conventional error-control techniques beyond the constraints of the Hamming metric by providing resilience against synchronization errors, particularly insertions and deletions. This enables more reliable information processing, transmission, and storage in developing high-density, high-throughput environments. The nature of the underlying noise fundamentally influences how function values are distorted during transmission. A function that is robust to substitution errors may behave unpredictably under insertions, deletions, erasures, or symbol-dependent noise, and the structural interactions between the function and the channel can vary dramatically. Developing FCCs tailored to specific channels enables reliable computation in settings where full message recovery is either unnecessary or prohibitively expensive, while also revealing deeper insights into how functions transform under diverse error processes. Such channel-aware FCC frameworks are essential for modern applications ranging from asynchronous communication and DNA storage to distributed sensing and edge computing, where computation must remain accurate despite highly nonuniform and context-dependent noise.
1.2 Related Works
@@
Function-correcting codes (FCCs) were introduced by Lenz et al. in [1] as a coding framework for recovering the value of a target function from the data rather than the data itself, resulting in substantially lower redundancy than traditional error-correcting codes. Their work formalized FCCs, established an equivalence with irregular-distance codes, and provided general upper and lower bounds demonstrating how the structure of the protected function influences the required redundancy. Subsequent developments, such as [6], revisited the model and extended the framework from the Hamming metric to the symbol-pair metric, deriving several bounds and structural insights. In [7], the authors established a lower bound on the redundancy of FCCs for linear functions. Focusing on the Hamming weight and Hamming weight distribution, Ge et al. [8] produced stronger lower bounds on redundancy and introduced code designs that exactly meet those bounds.
@@
A growing line of research has further generalized FCCs to nonstandard error models. In particular, [9] extended function-correcting codes to the $b$-symbol read channel, addressing clustered read errors common in modern storage systems. That work introduced irregular $b$-symbol distance codes and derived bounds characterizing the redundancy needed to recover a function value under this metric. The results, supported by a graph-theoretic formulation and illustrative examples, show that FCCs can achieve substantially lower redundancy than traditional $b$-symbol error-correcting codes when only the function output must be protected.
@@
Further extensions were provided in [10], which studied function-correcting codes for locally bounded $b$-symbol functions over $b$-symbol read channels. Recently, function correction has also been explored in the Lee-metric setting [11, 10]. More recent work, such as [12], investigates hybrid protection models in which both the function value and selected parts of the data must remain resilient to errors, highlighting the versatility of FCCs as a targeted and efficient alternative to full-message protection.
1.3 Contributions
@@
1.4 Organization
The rest of the paper is organized as follows. Section II provides a comprehensive review of the foundational concepts used throughout the paper. We begin by recalling the basic definitions and properties of insertion–deletion (insdel) codes, including the insdel metric and several equivalent formulations that are relevant to our analysis. In Section III, we introduce the framework of function-correcting deletion codes, function-correcting insertion codes, and function-correcting insertion–deletion (insdel) codes, which are designed to recover the value of a target function in the presence of deletion, insertion, or combined insertion–deletion errors. We then present and discuss irregular insdel-distance codes, highlighting their significance in characterizing optimal redundancy and in deriving key performance limits. Building on this framework, we derive lower and upper bounds on the optimal redundancy achievable by function-correcting codes for insdel channels. In addition, we establish Gilbert–Varshamov-type and Plotkin-like bounds on the lengths of irregular insdel-distance codes, which play a central role in analysing and understanding the redundancy of such codes. In Section IV, we apply the general theoretical results to a specific class of functions, the VT syndrome function, which plays a central role in the insertion–deletion coding. We analyze how the structure of the VT syndrome function interacts with insdel errors and derive both upper and lower bounds on the optimal redundancy required for function correction in this setting. We illustrate the general framework using three concrete classes of functions: the number-of-runs function (Section V), the maximum-run-length function (Section VI), and locally bounded functions (Section VII). For each class, we derive the corresponding redundancy values and their associated upper and lower bounds. Section VIII concludes the paper.
1.4.1 Notations
@@
2 Preliminaries
In this section, we review the basic concepts and definitions used throughout the paper. We begin with a brief discussion of insertion, deletion, and insdel codes, which are designed to enable reliable communication over channels that produce synchronization errors. We then introduce the fundamental notions associated with irregular-distance codes, a class of codes characterized by non-uniform or structure-dependent distance measures. Throughout the paper, we focus on codes over the alphabet $\mathbb{F}_{2}=\{0,1\}$.
2.1 Insdel code
In this paper, we consider a channel model in which errors occur as deletions, insertions, or a combination of both. Codes designed to handle such types of errors were first introduced by Levenshtein, Varshamov, and Tenengolts in the 1960s [2, 3, 4], and [5]. In this section, we provide a brief overview of these classical codes and summarize the main results associated with them.
Let $x$ be a binary vector of length $k$ that, when transmitted through a deletion channel, may lose some of its bits. If the channel deletes $t$ bits, the resulting binary vector $y$ has length ktk-t. The vector $y$ is referred to as a subsequence of $x$. A subsequence of a vector is obtained by selecting a subset of its symbols and aligning them in their original order, without any rearrangement. Formally, the subsequence is defined as follows:
Definition 2.1 (Subsequence).
@@ 2.1 Insdel code
Similarly, when $x$ is transmitted through an insertion channel, additional bits may be inserted at arbitrary positions, resulting in a supersequence of $x$. In the deletion case, the original vector $x$ is referred to as a supersequence of the received vector $y$. The set of all subsequence of $x$ obtained by deletion of $t$ bits is denoted by $D_{t}(x)$ and set of supersequences obtained by insertion of $t$ bits is denoted by $I_{t}(x)$.
@@
Definition 2.2 (Deletion Correcting Codes).
[3] A $t$-deletion-correcting code $\mathcal{C}$ of length $n$ is a subset of binary vector space $\mathbb{F}_{2}^{n}$ that holds the following property for all vectors $x,y\in\mathcal{C}$.
$D_{t}(x)\cap D_{t}(y)=\emptyset.$
Definition 2.3 (Insertion Correcting Codes).
[3] A $t$-insertion correcting code $\mathcal{C}$ of length $n$ is a subset of binary vector space $\mathbb{F}_{2}^{n}$ that holds the following property for all vectors $x,y\in\mathcal{C}$.
$I_{t}(x)\cap I_{t}(y)=\emptyset.$
@@
Definition 2.4 (Insdel Metric).
[3] Let $x$ and $y$ be binary sequences of length $k$. The insdel metric $d_{ID}(x,y)$ between $x$ and $y$ is defined as the minimum number of insertions and deletions required to transform $x$ into $y$.
@@
Definition 2.6.
(Insdel Code)[13] A subset of $\mathbb{F}^{k}_{2}$ is defined to be an insdel code $\mathcal{C}$ whose minimum distance is given by:
@@
Remark 2.7.
[3] Any code that can correct $s$ deletions (or equivalently, any code that can correct $s$ insertions) can also correct $s$ combined deletions and insertions.
@@
Lemma 2.8.
[14] Let $x,y$ be two binary words of length $k$. If $d_{H}(x,y)$ denotes the hamming distance between $x$ and $y$ then we have:
@@
Insertion and deletion errors can disrupt the alternating structure of a binary string. A useful combinatorial invariant for studying such transformations is the run count. The following lemma gives a lower bound on the insdel distance in terms of the difference in run counts. This bound will later help us establish limits on code correction and code construction capabilities.
Lemma 2.9.
@@
Assume without loss of generality that $r(y)\geq r(x)$. Let $m$ denote the number of insertions (and hence also the number of deletions) in an optimal insertion-deletion transformation from $x$ to $y$, so that the total number of operations is $d_{ID}(x,y)=2m$.
Let $A$ be the total change in run count caused by the $m$ insertions, and let $B$ be the total change in run count caused by the $m$ deletions. Then
r(y)r(x)=A+B.r(y)-r(x)=A+B.
Since a deletion can never increase the number of runs in a binary string, we have $B\leq 0$. Moreover, each insertion can increase the number of runs by at most $2$, and hence
$A\leq 2m.$
Combining these inequalities yields
$r(y)-r(x)=A+B\leq A\leq 2m.$
Therefore,
$m\geq\Big\lceil\frac{r(y)-r(x)}{2}\Big\rceil.$
Recalling that $d_{ID}(x,y)=2m$, we obtain
$d_{ID}(x,y)\geq 2\Big\lceil\frac{r(y)-r(x)}{2}\Big\rceil.$
If $r(x)\geq r(y)$, the same argument applies after swapping the roles of $x$ and $y$, yielding the bound in terms of $\lvert r(x)-r(y)\rvert$. Hence, in general,
$d_{ID}(x,y)\geq 2\Big\lceil\frac{\lvert r(x)-r(y)\rvert}{2}\Big\rceil.$
If $\lvert r(x)-r(y)\rvert$ is odd, then
$2\Big\lceil\frac{\Delta}{2}\Big\rceil=\Delta+1,$
where $\Delta=\lvert r(x)-r(y)\rvert$, completing the proof. ∎
We illustrate the tightness of the bound in Lemma 2.9 with two minimal examples, considering both even and odd run-count differences.
Example 2.10.
Even difference. Let $y=0100$ and $x=0000$. Here $r(y)=3$, $r(x)=1$, so |r(y)r(x)|=2|r(y)-r(x)|=2. The bound yields
$d_{ID}(x,y)\geq 2\Big\lceil\frac{2}{2}\Big\rceil=2.$
The insdel distance between $x$ and $y$ is given by
$d_{ID}(x,y)=2(4-3)=2.$
Hence, the bound is tight.
Odd difference. Let $y=0001$ and $x=0000$. Here $r(y)=2$, $r(x)=1$, so |r(y)r(x)|=1|r(y)-r(x)|=1. The bound yields
$d_{ID}(x,y)\geq 2\Big\lceil\frac{1}{2}\Big\rceil=2.$
The insdel distance between $x$ and $y$ is given by
$d_{ID}(x,y)=2(4-3)=2.$
Hence, the bound is tight in this case too.
For even number $2\leq d\leq 2k$, denote
@@
as the insdel ball centered at $x$ of radius $d$. Then, we have the following bounds on $|B_{ID}(x,d)|$.
Lemma 2.11.
@@
$|B_{ID}(x,2t)|\leq\binom{r(x)+t-1}{t}\Bigg(\sum_{i=0}^{t}\binom{k}{i}\Bigg)\leq\Bigg(\frac{e^{2}(k+t)k}{t^{2}}\Bigg)^{t}$
where $r(x)$ is the number of runs in $x$. Moreover, defining the worst-case insdel ball size as
$B^{\max}_{ID}(k,2t)\triangleq\max_{x\in\mathbb{F}_{2}^{k}}|B_{ID}(x,2t)|,$
we have
$B^{\max}_{ID}(k,2t)\leq\left(\frac{e^{2}(k+t)k}{t^{2}}\right)^{t}.$
Proof.
The first inequality in the upper bound follows from [15, Theorem 4]. For the second inequality, since the number of runs satisfies $r(x)\leq k$. Hence,
$\binom{r(x)+t-1}{t}\sum_{i=0}^{t}\binom{k}{i}\;\leq\;\binom{k+t-1}{t}\sum_{i=0}^{t}\binom{k}{i}$.
Using the standard estimates $\binom{a}{b}\leq\left(\tfrac{ea}{b}\right)^{b}$ and $\sum_{i=0}^{t}\binom{k}{i}\leq\left(\tfrac{ek}{t}\right)^{t}$, we obtain
$\binom{k+t-1}{t}\sum_{i=0}^{t}\binom{k}{i}\;\leq\;\left(\frac{e(k+t)}{t}\right)^{t}\left(\frac{ek}{t}\right)^{t}\;=\;\left(\frac{e^{2}(k+t)k}{t^{2}}\right)^{t}.$
Remark 2.12.
In the literature, an insdel ball is commonly referred to as a fixed-length Levenshtein ball. For a given center word, it is defined as the collection of all sequences of the same length that can be obtained from the center word by performing exactly $t$ deletions and exactly $t$ insertions. Equivalently, this set consists of all words whose insdel distance from the center word is at most $2t$ and whose length is preserved.
Lemma 2.13 (Singleton Bound).
[14] For a binary insdel code $\mathcal{C}$ of length $n$ and minimum distance $d$ , one has
$|\mathcal{C}|\leq 2^{n-d/2+1}$
Lemma 2.14.
@@
2.2 Function-Correcting Codes(FCC)
Definition 2.15 ([1]).